Summary
- Moved SDB1 along the Z axis by 3.6mm , and reduced the amount of DARM HF line beating in OMC reflection to zero
- Motion of SDB1 was done with BS TY loop open, and during the scan the CMRF has improved
- Turned back on the PSTAB1 line (~230Hz)
Figure 1 and 2 summarize the adjustment of SDB1 in Z that was made across two LN2 locks.
Figure 3 is a crude test with PSTAB loop open for 20s to check that the noise projection of PSTAB in the noise budget is working correctly, at least it does work above 100Hz.
Left interferometer in LN3, with MICH SET loop open
Notes from the shift
~13:25 UTC Turned off BS TY loop and put back BS TY INPUT to zero
13:34 UTC unlock
relock, B1p 12MHz mag 1-1.5e-3, with 7mW on 6MHz USB looking at the
peak before the carrier TEM00
disabled BS TY loop at the beginning of LN2
OB F0 X set at 1500
14:50 starting to decrease OB F0 X
unlocked with OB F0 X at 800
adjusted SDB1 TX set point from 43 to 59 and put to fixed set points,
so the bench is at the position that was aligned in LN2 according to
OMC AA signals
turned back on PSTAB1 line
unlock at LN1 transition
found issue with following two unlocks at LN1. Putting back floating
set point did not work, and it was still on fixed set points that were
6urad away, and jumping to those during the LN1 transition. Pressing
the floating set point button again solved the issue.
Adjusted MICH offset to reduce B4 56MHz sideband imbalance and locked
in LN2, opened the BS TY loop
moving SDB1 in Z, and adjusting in TX at the same time (2urad per
100um on OB F0 X, 150s long steps)
Arrived at zero of SDB1_B1s_DARM_HF_i around 20:13 UTC. Did not try to
go further to see if more improvements are possible there. OB F0 X set at -2100.
~20:24 UTC Turning off MICH SET loop, and adjusting it to have small B1p 12MHz
and balanced B4 56Mhz sidebands. And turning back on BS TY loop.
20:52:20 (20s) opened PSTAB loop, but immediately SR started moving in
alignment by 0.2urad, closed it back up